OpenStudy (anonymous):

Solve using the multiplication principle -6x>1/13

OpenStudy (anonymous):

x|x>2 1/6

OpenStudy (saifoo.khan):

\[x < \frac{1}{78}\]

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

note: when multiplying/dividing by -1 the inequality is reversed,

OpenStudy (anonymous):

so it is just 78

OpenStudy (saifoo.khan):

No.

OpenStudy (saifoo.khan):

x < 1/78

OpenStudy (anonymous):

ok

OpenStudy (unklerhaukus):

saifoo has solved the correctly

OpenStudy (unklerhaukus):

this*

OpenStudy (unklerhaukus):

*except he forgot to keep the minus sign -6x>1/13 =-x>1/78 =x<1/-78